# An Empirical Study of Self-Admitted Technical Debt in Machine Learning Software

### v2

@@ 6.2. Construct Validity
There is no consensus among researchers regarding the inclusion of missed requirements or known defects in the concept of SATD. The seminal work by Kruchten et al.(kruchten2013technical, ) does not regard requirement and defect debt as SATD, however, the work of Bavota et al. (bavota2016large, ) does consider the two categories as SATD. Despite these differences, as the majority of research ((alves2014towards, ; maldonado2015detecting, ; bavota2016large, ; da2017using, ; wattanakriengkrai2018identifying, ; liu2020using, )) on SATD accepts Requirement and Defect Debt as SATD, we include these categories in our manual analysis for RQ2. This potential variation in the understanding and interpretation of SATD poses a construct validity threat as it could lead to varying interpretations and thus results.
The curation of ML repositories involved two levels, with an initial level using programmatic checks to verify criteria C1, C2, and C3, and with level-2 using manual checks against criteria C4 and C5. Our Level-2 checks were done by one author, which represents a minor threat to validity due to subjective bias. We believe that the threat is minimal since for both ML and non-ML repositories, the first step is to check the project’s Readme itself for the criteria C4 and C5, which is a straightforward process. Typically, README files in software development include sections on the project’s purpose, installation instructions, usage examples, etc. (prana2019categorizing, ). While a programmatic string-matching-based tool could potentially scan for ML-related keywords (eliminating the need for a manual analysis altogether), our manual efforts were only done to make the process more resilient to errors. Furthermore, checking for criteria C4 and C5 using a highly structured approach as outlined in sub-section 4.1.3 is a curation process, inherently objective in nature, which can minimize subjective bias (hallowell2009techniques, ). This manual curation for repository selection is opposed to the qualitative analysis done for RQ2 and RQ3, where more substantial qualitative coding is required, for which we performed multi-rater coding with high inter-rater agreements.
@@ 6.3. External Validity
Finally, the results of our findings are dependent on our sampled data. While the sample of ML projects was based on exhaustive search (i.e., including relevant projects in a domain until subsequent projects are irrelevant), procuring non-ML projects was based on convenience sampling (getting the top-318 non-ML projects as long as the selection criteria are met). Although a random sample would offer a more unbiased comparison, obtaining all non-ML Python projects across all GitHub pages, then selecting a random sample may not be feasible.
Moreover, although we select mature projects by using selection criteria inspired by related work (kalliamvakou2014promises, ; chen2020studying, ; alfadel2021use, ) such as a project requiring more than a month of development history, it is still possible that immature projects may seep into our analysis, biasing our results. Following prior research (alfadel2021use, ; chen2020studying, ) that has commonly mitigated this threat using threshold-based criteria, we establish five filtering criteria (e.g., a project should have more than one month of developmental history) for the removal of toy projects. Notably, our ML dataset lacks test cases. This absence signifies a deviation from traditional Software Engineering where tests involve unit tests and integration tests; and code quality is measured through the coverage of such tests. However, such testing practices do not seamlessly translate to ML scenarios (zhang2020machine, ) [^21] Accessed on: 2024-03-22, https://fg-tav.gi.de/fileadmin/FG/TAV/47.TAV/TAV47-Felderer-TestingTheUntestable.pdf [^22] Accessed on: 2024-03-22, https://shorturl.at/dnBDU. Instead of test cases, ML testing often involves ensuring that a model has high performance, high robustness, lack of bias, etc. Indeed, the lack of quality assurance standards (for instance % test coverage which is easily measured in traditional software) in ML systems, in itself is a motivation for our study.
We conducted an analysis of a diverse portfolio consisting of 318 projects from well-established ML domains like “self-driving car” and “natural language processing” to ensure that our findings are representative of industry-grade ML projects. However, our results may not encapsulate the SATD phenomena across the entire spectrum of ML software systems.
@@
In RQ3, our analysis was confined to the examination of SATD within five stages (i.e., Model Building, Data Prepossessing, Data Reading, Model Validation, and Model Deployment) of the ML pipeline as proposed by Amershi et al. (amershi2019software, ). Consequently, the identified technical debt may not fully represent the nuances of TD that could occur in expanded areas of ML pipelines like model monitoring, maintenance, and evolution. Future studies could aim to analyze TD in these expanded areas or in MLOps development and monitoring tasks to provide a more comprehensive understanding of TD in overall ML-based software development. This limitation suggests that our results are primarily applicable to the specific context we investigated, and caution should be taken when generalizing the findings to these other contexts.
@@ 6.4. Conclusion Validity
Just like debt in monetary situations, it may be argued that not all technical debt is bad and needs to be paid off eventually. Technical debt (TD) manifests in various forms, such as organizational, discussion, and product-related TD, and not all TD is detectable by tools. Prior studies by Potdar and Shihab (potdar2014exploratory, ) recognized SATD as an effective heuristic for estimating TD in software systems. The importance of understanding SATD for software maintenance and management has been acknowledged in previous non-ML research (Li2015, ; Alves2016, ; Rios2018, ; Alfayez2020, ; Lim2012, ; wehaibi2016examining, ; brown2010managing, ; letouzey2012managing, ; kruchten2012technical, ; kruchten2013technical, ; bavota2016large, ), and generally the prevalence of SATD is considered a bad code smell (fowler2018refactoring, ).
@@
7. Implications
Implications for ML Practitioners: ML practitioners should be more vigilant in providing better documentation for technical debt and be aware of the quick accumulation of debt. The results from RQ1 demonstrate that ML code contains a significantly higher proportion of technical debt compared to non-ML code, confirming the observation that ML code is highly debt-prone. The findings from RQ2 indicate that a substantial portion of this debt is related to requirements, suggesting that ML engineers should acknowledge sub-optimal implementations or missing features, planning for their implementation in the future. To address this, ML practitioners should strive to improve communication and collaboration between ML, non-ML, and data developers ensuring that technical debt is managed effectively across the entire software system. Leveraging tools (like DVC, MLflow) and adopting best practices that facilitate this collaboration can be beneficial.
Furthermore, our newly-found SATD type, inadequate testing indicates the presence and need to improve ML testing to increase the reliability of code functionalities. Untested pieces of code could result in hidden ML bugs (due to the stochastic nature of ML) or manifest as sub-optimal ML model performance.
RQ5 reveals that technical debt introduced during large code changes spanning multiple low-complexity files tends to linger longer in the source code. ML practitioners should pay special attention to such commits during code review activities and quickly refactor the corresponding technical debts when the code complexity is still low. While our results are based on our ML dataset, future research may wish to explore the generalizability of our findings on non-ML data as well.
In RQ4, we observed that ML code quickly accumulates debt, however debt removal is also quick. We attribute such quick accumulation and removal to a high ML churn, as shown in RQ1. ML developers engaged in such high churn software development should be cautious by prioritizing code reviews, automating testing, and ensuring proper documentation. Given the highly experimental nature of ML pipelines, developers may also wish to employ the use of feature flags, to turn on and off specific functionalities to allow for safe experimentation without constantly altering the codebase.
Overall, ML engineering teams should also focus on improving code documentation, implementing effective test cases, and ensuring the reliability of ML pipelines and models. By enforcing abstraction barriers, increasing code readability and documentation, avoiding unnecessary dependencies and reducing complexity, ML practitioners can improve the overall software quality and maintainability of their ML systems. Frequent refactoring and restructuring of software system design would help remove technical debt and reduce complexity.
Implications for Researchers and Toolsmiths: The findings from RQ1, RQ2, and RQ3 highlight the prevalence of technical debt in ML code, particularly in relation to requirements, testing, and configuration. Researchers and toolsmiths should develop techniques and tools that support ML engineering teams in efficiently handling requirements, managing data dependencies and model dependencies effectively, and improving testing infrastructure and configuration management. There is a need for tools that detect and remove technical debt, estimate repayment efforts, and assess the impact of different types of technical debt to prioritize their resolution during maintenance activities. Researchers should also investigate the implications of technical debt in emerging ML paradigms, such as federated learning and reinforcement learning, and develop tailored strategies for addressing technical debt in these contexts. Section 8 provides an in-depth explanation of the potential research and tooling avenues inspired by our research.
@@ 7. Implications
Implications for ML Educators: Given the prevalence of technical debt in ML code, ML educators should incorporate the concepts of technical debt management, software engineering best practices, and documentation strategies in their ML courses and training programs. By educating the next generation of ML practitioners about the potential pitfalls and long-term consequences of technical debt, ML educators can contribute to improving the overall quality and maintainability of ML systems. Case studies and real-world examples should be included to demonstrate the impact of technical debt on ML projects, fostering a deep understanding of the consequences of neglecting technical debt. Given that requirement debt is most prevalent (RQ2), and the model building side accumulates the maximum amount of debt (RQ3), ML educators may wish to provide special focus on such areas. This approach can help ML practitioners to be better equipped to manage technical debt effectively.
@@ 8. Future Work
  • Examining TD (SATD) in Other Aspects of ML Development. In RQ3, we label 611 ML Self-Admitted Technical Debt (SATD) instances onto the ML components proposed by Amershi et al (amershi2019software, ). Future research may consider examining Technical Debt (TD) in ML components that are not mentioned in Amershi’s pipeline. This could involve supporting tasks like model monitoring, maintenance, and evolution, ethics and fairness evaluation, or the user feedback loop. Evaluating SATD in ML continuous integration/ continuous delivery (CI/CD) tasks like model/data versioning, EOL (end of life planning), or rollback and redeployments, etc. could also be a valuable area for the research community. Overall, future work could investigate TD’s impact on the performance and robustness of ML models, as well as the overall system’s reliability, security, and maintainability.
@@
  • Leveraging ML for Detecting ML-specific SATD. Existing SATD detection tools mainly target non-ML software. Due to the distinct nature of ML, there might be a need for ML-specific SATD detection tools that are more effective in identifying ML-related debt. Developing such tools in future research could facilitate efficient identification, tracking, and resolution of SATD throughout ML development, covering a wide range of ML tasks (amershi2019software, ).
@@
  • Evaluation of SATD Pre-Rebasing. Our research primarily examines the spread of SATD within the master branch. However, future studies might consider investigating the behavior of SATD across various branches. Given that change evaluation across branches is a burgeoning field of research (bhatia2023towards, ), examining SATD dynamics in this context could be valuable in understanding debt.
@@ 9. Conclusion
For RQ1, we find that Machine Learning (ML) code accumulates twice the debt in comparison to non-ML code. While Sculley et al. (Sculley, ) previously hypothesized the potential for debt to cause maintenance challenges in ML software, ours is the first study to provide empirical evidence for the pronounced prevalence of debt in ML code. Notably, our finding that 2% of non-ML code has SATD augments findings from prior SATD studies on non-ML software. Our results align and further contextualize the research by Potdar and Shihab (potdar2014exploratory, ), who ascertained that debt permeates in 2.7% to 31% of files, 0.4% to 3.3% at a class level, and 0.3% to 2.6% at a function level. Our methodology quantifies debt based on the percentage of comments marked as SATD, facilitating a comparative analysis between ML and non-ML comments.
@@
In RQ2, our qualitative efforts extend the prior taxonomy of SATD. The prior taxonomy, by Bavota and Russo (bavota2016large, ), proposed seven years ago, necessitates augmentation to cater to the ML-dominated software development landscape that is more relevant today. In our qualitative analysis of SATD in ML code, we identified two new ML-specific SATD categories: “inadequate testing” and “configuration” debts. We believe that these categories emerged due to the unique challenges of ML pipeline development, which involves frequent experimentation for model optimization and management of hyperparameters and data preprocessing configurations. These findings offer a clearer view of the technical debts present in today’s ML-centric software development.
@@
In RQ3, we analyzed SATD within different stages of Amershi et al.’s ML pipeline (amershi2019software, ). This not only categorizes the debt but also pinpoints the stages in the ML pipeline where technical debt is more likely to occur, offering a structured approach to address ML debts.
@@
For RQ4, we examined how quickly ML and non-ML debts are introduced and subsequently resolved. Our data shows that ML debts are introduced by 140 times and removed by 3.7 times more rapidly than non-ML debts. This aligns with the insights from Sculley et al. (Sculley, ), which underscore the critical nature of proficient debt management in ML development. Our results complement existing research on SATD survival in non-ML software. Bavota and Russo (bavota2016large, ) reported debt survival in terms of commits and reported that SATD survives for a median of 266 commits, yet did not provide time-related results. Conversely, Maldonado et al. (maldonado2017empirical, ) indicated that SATD survives between 18.2-172.8 days, averaging at 82-613.2 days. While Maldonado et al. presented average timelines for debt survival, we employed survival analysis using the Kaplan-Meier metric. This approach offers more robust insights than raw averages because it accounts for censored data, providing a comprehensive view of life expectancy of SATD.
@@
In RQ5, using a ML model interpretation, we delve into the characteristics of long-lasting debts in ML code. While past research briefly touched upon debt duration, to our knowledge, our study is the first to uncover an understanding of long-lasting debts, setting it apart from both ML and non-ML research in the area. Moreover, we use software metrics in the context of SATD. While earlier studies such as (yan2018automating, ) utilized software metrics to detect SATD using ML models, our approach leverages these metrics to interpret the specific traits of long-lasting debt. Our findings indicate that files with lower complexity are more likely to permeate debt until a long time, suggesting developers should give heightened attention to such potentially overlooked small files.
@@
These results provide insights to practitioners towards better management of technical debt and call for more research work to improve our understanding of technical debt occurring in ML-based systems.

### v3

@@ 6.2. Construct Validity
There is no consensus among researchers regarding the inclusion of missed requirements or known defects in the concept of SATD. The seminal work by Kruchten et al.(kruchten2013technical) does not regard requirement and defect debt as SATD, however, the work of Bavota et al. (bavota2016large) does consider the two categories as SATD. Despite these differences, as the majority of research ((alves2014towards; maldonado2015detecting; bavota2016large; da2017using; wattanakriengkrai2018identifying; liu2020using)) on SATD accepts Requirement and Defect Debt as SATD, we include these categories in our manual analysis for RQ2. This potential variation in the understanding and interpretation of SATD poses a construct validity threat as it could lead to varying interpretations and thus results.
The curation of ML repositories involved two levels, with an initial level using programmatic checks to verify criteria C1, C2, and C3, and with level-2 using manual checks against criteria C4 and C5. Our Level-2 checks were done by one author, which represents a minor threat to validity due to subjective bias. We believe that the threat is minimal since for both ML and non-ML repositories, the first step is to check the project’s Readme itself for the criteria C4 and C5, which is a straightforward process. Typically, README files in software development include sections on the project’s purpose, installation instructions, usage examples, etc. (prana2019categorizing). While a programmatic string-matching-based tool could potentially scan for ML-related keywords (eliminating the need for a manual analysis altogether), our manual efforts were only done to make the process more resilient to errors. Furthermore, checking for criteria C4 and C5 using a highly structured approach as outlined in sub-section 4.1.3 is a curation process, inherently objective in nature, which can minimize subjective bias (hallowell2009techniques). This manual curation for repository selection is opposed to the qualitative analysis done for RQ2 and RQ3, where more substantial qualitative coding is required, for which we performed multi-rater coding with high inter-rater agreements.
@@ 6.3. External Validity
The results of our findings are dependent on our sampled data. To compare ML and Non-ML Projects in RQ1 and RQ4, we have taken steps to ensure that the selection process is consistent for both ML and Non-ML fronts. While the sample of ML projects was based on exhaustive search (i.e., including relevant projects in a domain until subsequent projects are irrelevant), procuring non-ML projects was based on convenience sampling (getting the top-318 non-ML projects as long as the selection criteria are met). Although a random sample would offer a more unbiased comparison, obtaining all non-ML Python projects across all GitHub pages, then selecting a random sample may not be feasible.
Moreover, although we select mature projects by using selection criteria inspired by related work (kalliamvakou2014promises; chen2020studying; alfadel2021use) such as a project requiring more than a month of development history, it is still possible that immature projects may seep into our analysis, biasing our results. Following prior research (alfadel2021use; chen2020studying) that has commonly mitigated this threat using threshold-based criteria, we establish five filtering criteria (e.g., a project should have more than one month of developmental history) for the removal of toy projects. Notably, our ML dataset lacks test cases. This absence signifies a deviation from traditional Software Engineering where tests involve unit tests and integration tests; and code quality is measured through the coverage of such tests. However, such testing practices do not seamlessly translate to ML scenarios (zhang2020machine) [^23] Accessed on: 2024-03-22, https://fg-tav.gi.de/fileadmin/FG/TAV/47.TAV/TAV47-Felderer-TestingTheUntestable.pdf [^24] Accessed on: 2024-03-22, https://shorturl.at/dnBDU. Instead of test cases, ML testing often involves ensuring that a model has high performance, high robustness, lack of bias, etc. Indeed, the lack of quality assurance standards (for instance % test coverage which is easily measured in traditional software) in ML systems, in itself is a motivation for our study.
We conducted an analysis of a diverse portfolio consisting of 318 projects from well-established ML domains like “self driving car” and “natural language processing” to ensure that our findings are representative of industry-grade ML projects. However, our results may not encapsulate the SATD phenomena across the entire spectrum of ML software systems.
@@
In RQ3, our analysis was confined to the examination of SATD within five stages (i.e., Model Building, Data Prepossessing, Data Reading, Model Validation, and Model Deployment) of the ML pipeline as proposed by Amershi et al. (amershi2019software). Consequently, the identified technical debt may not fully represent the nuances of TD that could occur in expanded areas of ML pipelines like model monitoring, maintenance, and evolution. Future studies could aim to analyze TD in these expanded areas or in MLOps development and monitoring tasks to provide a more comprehensive understanding of TD in overall ML-based software development. This limitation suggests that our results are primarily applicable to the specific context we investigated, and caution should be taken when generalizing the findings to these other contexts.
@@ 6.4. Conclusion Validity
Just like debt in monetary situations, it may be argued that not all technical debt is bad and needs to be paid off eventually. Technical debt (TD) manifests in various forms, such as organizational, discussion, and product-related TD, and not all TD is detectable by tools. Prior studies by Potdar and Shihab (potdar2014exploratory) recognized SATD as an effective heuristic for estimating TD in software systems. The importance of understanding SATD for software maintenance and management has been acknowledged in previous non-ML research (Li2015; Alves2016; Rios2018; Alfayez2020; Lim2012; wehaibi2016examining; brown2010managing; letouzey2012managing; kruchten2012technical; kruchten2013technical; bavota2016large), and generally the prevalence of SATD is considered a bad code smell (fowler2018refactoring).
@@
7. Implications and Recommendations for ML Stakeholders
Our RQ1 evidence shows that ML repositories accumulate SATD twice as fast as comparable non-ML code. Debt clusters in two main hotspots: model-building (47%) and data-pre-processing ($\approx$ 25%), and are dominated by functional-requirement debt (34%), configuration debt (12%), and the newly observed inadequate testing debt (5%). Based on these specific patterns, we recommend implications for practitioners, researchers, toolsmiths, and educators.
1. Address configuration debt early For ML software, configuration debt is one of the prominent ML-type debt. Some prominent examples from our analysis include “# hardcoded search strategy”252525last accessed Apr 17, 2025 https://github.com/chris-chris/pysc2-examples/blob/master/a2c/kfac.py, or “# TODO test with different speeds”262626last accessed Apr 17, 2025 https://github.com/NVIDIA/DeepLearningExamples/blob/master/CUDA-Optimized/FastSpeech/generate.py or “# TBD: use better initializers (uniform, etc.)”272727last accessed Apr 17, 2025 https://github.com/kermitt2/delft/blob/master/delft/sequenceLabelling/preprocess.py show notable examples of configuration debt that is likely to accumulate in ML systems. An actionable item for ML practitioners would be to remain vigilant towards the accumulation of configuration debt by ensuring that all hyperparameter settings and experimental configurations are systematically tracked. This can be achieved through the use of experiment-tracking tools such as DVC or MLflow, where configurations are stored in version-controlled YAML or JSON files rather than being hardcoded within the source code (sklavenitis2025scoping). To enforce this, pre-commit hooks or continuous integration (CI) pipelines can be employed to automatically reject code submissions that introduce ad-hoc constants or fail to reference external configuration files. Moreover, each configuration file can be tagged with an assigned owner responsible for its maintenance, along with a predefined review date or expiry period. Upon reaching this date, CI systems can be configured to automatically generate review tasks or issues, prompting the team to reassess the relevance and accuracy of the configuration. Such practices not only minimize the accumulation of configuration debt but also enhance reproducibility and maintainability within ML systems (sklavenitis2025scoping). Toolsmiths can build CI tools that have ability to enforce such settings.
2. Align Technical Debt Remediation times (budget payment windows) with High-Churn Development Phases Our survival analysis (RQ4) shows that SATD in ML code typically appears within 10 days of file creation, compared to 1,400 days in non-ML code. Additionally, our findings (RQ5) indicate that SATD often arises during commits that modify many files but exhibit low per-file complexity, suggesting rapid, large-scale changes with limited design consideration. These patterns point to quick, exploratory development cycles as a key contributor to persistent technical debt.
To mitigate this, we recommend that ML practitioners should proactively schedule targeted refactoring shortly after such high-churn activities. Version control systems or CI pipelines can specifically monitor the ratio of normalized churn (i.e., churn to lines of code), teams can identify commits in the top quartile of normalized churn, and flag such commits for careful review/refactoring. Later in the budget friendly periods of downtime where delivery pressure is manageable, debt can be repaid with a special focus on high churn files, and those that were changed in alarming commits could be reviewed for potential SATD and considered for refactoring in the immediate following sprint. This approach aligns with our empirical observation that timely intervention can reduce the likelihood of SATD persisting long-term by approximately 20%. These strategies can systematically keep technical debt in check for the long-run, improving maintainability without requiring disruptive, large-scale overhauls.
3. Prioritize Rigorous Testing for Model-Building Components. Our study identifies a category of technical debt more-likely found in ML systems: inadequate testing debt. These are developer comments acknowledging missing or insufficient tests in critical parts of the ML pipeline, especially in model-building code (e.g., “TODO 1 Test the fully_connected and conv2d function”). Such gaps can lead to silent failures or degraded model performance, as untested ML components may behave unpredictably under certain inputs or could lead to ill-defined configurations, causing unwanted behaviours (like low accuracy) in models or leading to unstable model training (de2025classification; sklavenitis2024measuring).
To address this, we recommend that ML practitioners treat model-building code with the same testing rigor as traditional production code. Concretely, ML practitioners and Dev(ML)-ops toolsmiths can:
  • Introduce property-based or metamorphic tests to validate key invariants, for instance, consistent output shapes across inputs and reproducibility under fixed random seeds.
  • Enforce a minimum (say 80%) statement coverage for data preprocessing and model code before allowing merges. Such checks can be integrated in CI pipelines.
  • Implement regular drift testing: periodically (e.g., nightly), re-run the model on a frozen validation dataset and compare key performance metrics (e.g., accuracy, loss) against historical values. Triggered alerts can then be attributed to test coverages across Data and Model components, along with high-churn development phases, as mentioned previously.
This structured approach ensures that untested or poorly tested model logic is minimized, improving both code reliability and model performance stability.
4. Isolate experiments behind feature flags and short-lived branches. Our results for RQ1 and RQ4 indicate that, in ML systems, high churn and exploratory work is inevitable. It is possible that half implemented ideas (as indicated by code contaminated with large amounts of SATD) might seep into the main codebase by junior developers or during a hurried development sprint. Safeguarding against such experimental features seeping into the main developmental branch is essential. Practitioners can use runtime flags (for instance, Hydra [^28] Last accessed Apr 27, 2025 https://hydra.cc/docs/advanced/hydra-command-line-flags/ or LaunchDarkly292929last accessed Apr 17, 2025 https://launchdarkly.com) so researchers and practitioners can ship prototypes without altering the production path. Furthermore, Dev(or ML)-ops practitioners can enable CI rigs such that branches older than ‘N’ weeks without merges are auto-archived; flags older than one release trigger a “clean-up” review. This fortifies the gates to confine debt into explicitly experimental gated areas, and prevents accidental seepage into the main developmental branch. Empirical studies of feature toggle practices (mahdavi2021software) emphasize the need for systematic cleanup routines and lifecycle management, corroborating with our call to prevent long-lived flags from becoming a persistent source of technical debt.
5. Know how much you debt owe, and don’t suffer later: Implement debt heat-maps within the pull-request review workflow. Our analysis of debt survival indicates that debt can stay for a long time in both traditional and ML software. Technical debt can be a neglected area, and just like financial debt, not knowing how much you owe can be catastrophic to systems health. We recommend implementing SATD heat-maps visualizations within CI systems such as Jenkins303030last accessed Apr 17, 2025 https://www.jenkins.io. This heat-map can visualize the output of an automated SATD detection tool such as the one we used in our research (liu2018satd). By visualizing the code diff with intuitive colour-coded overlays, reviewers are immediately alerted to potential debt-prone regions—for example, large green-highlighted sections indicating extensive code modifications in low-complexity areas where SATD has been flagged. This visual augmentation of pull requests can also draw attention to structural weak points that might otherwise escape scrutiny. Even though payment of debt is done in the future, at least the debt-riddled commits and LOCs are not “forgotten” in the long run. Recently, tools like DebViz (li2023debtviz) provision identifying, measuring, visualizing, and monitoring SATD, answering to this call.
Moreover, setting Service-Level Objectives (SLOs) (hidalgo2020decision) for debt comments can help in the long run. For instance, adding a bot that pings the author if an SATD comment is still present two weeks after the introduction or past a release can prevent long-term accumulation of debt. Public dashboards showing open-vs-resolved counts might also create gentle social pressure and keep removal latency low.
@@ 7. Implications and Recommendations for ML Stakeholders
6. Integrate “debt labs” into ML curricula. By educating the next generation of ML practitioners about the potential pitfalls and long-term consequences of technical debt, ML educators can contribute to improving the overall quality and maintainability of ML systems. This implication aligns with several pedagogical studies (tonin2017effects; crespo2021carrot) focused on reducing technical debt among computer science students. Case studies and real-world examples (for instance, students must (i) identify SATD in an open-source ML repository, (ii) write a remediation plan, and (iii) implement one fix. This practice grounds theory in visible, real-world debt and builds habits that transfer to industry) should be included to demonstrate the impact of technical debt on ML projects, fostering a deep understanding of the consequences of neglecting technical debt. Given that requirement debt is most prevalent (RQ2), and the model building side accumulates the maximum amount of debt (RQ3), ML educators may wish to provide special focus on such areas. This approach can help the next generation of ML practitioners to be better equipped to manage technical debt effectively.
@@ 8. Future Work
  • Examining TD (SATD) in Other Aspects of ML Development. In RQ3, we label 611 ML Self-Admitted Technical Debt (SATD) instances onto the ML components proposed by Amershi et al (amershi2019software). Future research may consider examining Technical Debt (TD) in ML components that are not mentioned in Amershi’s pipeline. This could involve supporting tasks like model monitoring, maintenance, and evolution, ethics and fairness evaluation, or the user feedback loop. Evaluating SATD in ML continuous integration/ continuous delivery (CI/CD) tasks like model/data versioning, EOL (end of life planning), or rollback and redeployments, etc. could also be a valuable area for the research community. Overall, future work could investigate TD’s impact on the performance and robustness of ML models, as well as the overall system’s reliability, security, and maintainability.
@@
  • Leveraging ML for Detecting ML-specific SATD. Existing SATD detection tools mainly target non-ML software. Due to the distinct nature of ML, there might be a need for ML-specific SATD detection tools that are more effective in identifying ML-related debt. Developing such tools in future research could facilitate efficient identification, tracking, and resolution of SATD throughout ML development, covering a wide range of ML tasks (amershi2019software).
@@
  • Evaluation of SATD Pre-Rebasing. Our research primarily examines the spread of SATD within the master branch. However, future studies might consider investigating the behavior of SATD across various branches. Given that change evaluation across branches is a burgeoning field of research (bhatia2023towards), examining SATD dynamics in this context could be valuable in understanding debt.
@@ 9. Conclusion
For RQ1, we find that Machine Learning (ML) code accumulates twice the debt in comparison to non-ML code. While Sculley et al. (Sculley) previously hypothesized the potential for debt to cause maintenance challenges in ML software, ours is the first study to provide empirical evidence for the pronounced prevalence of debt in ML code. Notably, our finding that 2% of non-ML code has SATD augments findings from prior SATD studies on non-ML software. Our results align and further contextualize the research by Potdar and Shihab (potdar2014exploratory), who ascertained that debt permeates in 2.7% to 31% of files, 0.4% to 3.3% at a class level, and 0.3% to 2.6% at a function level. Our methodology quantifies debt based on the percentage of comments marked as SATD, facilitating a comparative analysis between ML and non-ML comments.
@@
In RQ2, our qualitative efforts extend the prior taxonomy of SATD. The prior taxonomy, by Bavota and Russo (bavota2016large), proposed seven years ago, necessitates augmentation to cater to the ML-dominated software development landscape that is more relevant today. In our qualitative analysis of SATD in ML code, we identified two new ML-specific SATD categories: “inadequate testing” and “configuration” debts. We believe that these categories emerged due to the unique challenges of ML pipeline development, which involves frequent experimentation for model optimization and management of hyperparameters and data preprocessing configurations. These findings offer a clearer view of the technical debts present in today’s ML-centric software development.
@@
In RQ3, we analyzed SATD within different stages of Amershi et al.’s ML pipeline (amershi2019software). This not only categorizes the debt but also pinpoints the stages in the ML pipeline where technical debt is more likely to occur, offering a structured approach to address ML debts.
@@
For RQ4, we examined how quickly ML and non-ML debts are introduced and subsequently resolved. Our data shows that ML debts are introduced by 140 times and removed by 3.7 times more rapidly than non-ML debts. This aligns with the insights from Sculley et al. (Sculley), which underscore the critical nature of proficient debt management in ML development. Our results complement existing research on SATD survival in non-ML software. Bavota and Russo (bavota2016large) reported debt survival in terms of commits and reported that SATD survives for a median of 266 commits, yet did not provide time-related results. Conversely, Maldonado et al. (maldonado2017empirical) indicated that SATD survives between 18.2-172.8 days, averaging at 82-613.2 days. While Maldonado et al. presented average timelines for debt survival, we employed survival analysis using the Kaplan-Meier metric. This approach offers more robust insights than raw averages because it accounts for censored data, providing a comprehensive view of life expectancy of SATD.
@@
In RQ5, using a ML model interpretation, we delve into the characteristics of long-lasting debts in ML code. While past research briefly touched upon debt duration, to our knowledge, our study is the first to uncover an understanding of long-lasting debts, setting it apart from both ML and non-ML research in the area. Moreover, we use software metrics in the context of SATD. While earlier studies such as (yan2018automating) utilized software metrics to detect SATD using ML models, our approach leverages these metrics to interpret the specific traits of long-lasting debt. Our findings indicate that files with lower complexity are more likely to permeate debt until a long time, suggesting developers should give heightened attention to such potentially overlooked small files.
Taken together, our results are better explained by development churn—rapid, iterative changes to models, datasets, and configurations that lead to frequent SATD introduction followed by prompt cleanup. This lens reconciles the faster addition/removal observed in RQ4 with the higher overall prevalence in RQ1: high-velocity iteration inflate short-lived debt while removals keep pace. It also contextualizes RQ5: the debts that persist despite churn tend to reside in low-complexity files.
@@
Our results provide insights to practitioners towards better management of technical debt and call for more research work to improve our understanding of technical debt occurring in ML-based systems.